OpenStudy (anonymous):

is the following function even, odd, or neither y = 7x^2 - 7x + 1

OpenStudy (campbell_st):

substitute -x for x and see what you get is f(-x) = f(x) its even f(-x) = - f(x) its odd or if f(-x) = neither of the above then its neither off nor even

OpenStudy (anonymous):

so its even

OpenStudy (campbell_st):

well f(-x) = 7(-x)^2 - 7(-x) + 1 = 7x^2 + 7x + 1 so its neither -f(x) = -7x^2 + 7x - 1 would make it odd and and its not even because you didn't get the original function

OpenStudy (anonymous):

RIGHT ITS NOT EVEN GOOD